# Disorder-induced strong-field strong-localization in 2D systems

### v1

@@ I Introduction
First, the system generically displays a hexagonal WC with explicit breaking of the rotational symmetry over a broad range of fillings, and an incompressible FQH liquid appears only in a very narrow window around the odd-denominator fraction such as $\nu=1/3$ Laughlin state (e.g., $\nu\approx 0.334$). Even a slight deviation to $\nu\approx 0.311$ or $0.356$ rapidly restores the WC pattern, consistent with the idea that the $\nu=1/3$ phase is stabilized only at (or extremely close to) commensurate odd-denominator fillings. The FQH state slightly lowers its energy compared with the background of a smooth WC phase (existing for all values of $\nu$) through small cusps at these precise odd-denominator $\nu$ (e.g., $1/3$ real space profile as can be seen clearly in Fig. 1 of Ref. 1). Although this is not theoretically unexpected [2, 3, 4], the actual observation of a generic WC at all $\nu$ around (within 5%) $\nu=1/3$ is really the first direct verification of this theoretical expectation of the FQHL states being stabilized (with respect to WC) only at precise odd denominator fillings. This experiment is also the first direct observation of the generic existence of the WC phase in the lowest Landau level of 2D systems.
A new experimental finding in Ref. 1, however, is that, as $\nu$ continues decreasing, eventually at a small $\nu=0.093\approx 1/11$, the electrons cross over to a completely spatially random amorphous solid phase. In the WC, the structure factor $S(q)$ exhibits six well-developed Bragg peaks, reflecting the long-range hexagonal order of the crystal, whereas in the FQHL $S(q)$ is essentially a featureless blob, consistent with the absence of long-range spatial correlations in real space. By contrast, in the amorphous solid phase the real-space images show electrons localized at random, quenched positions, and the corresponding $S(q)$ displays a characteristic hybrid structure: a diffuse central blob accompanied by a surrounding ring. The presence of the ring in $S(q)$ unambiguously identifies the amorphous solid as a solid with a well-defined characteristic length scale, while the central blob signifies liquid-like, short-range spatial correlations. Together, these features establish the amorphous solid as a distinct phase, solid in the sense of electron localization, yet lacking the long-range translational and rotational order characteristic of a crystalline WC. Another impressive recent zero-field density-tuned STM study has reported the observation of a low-density zero-field electronic amorphous phase in highly disordered bilayer MoSe2 samples (no moiré) [5], which has been interpreted as the manifestation of the strongly-localized random disorder-dominated zero-field AS phase in Ref. [6].
@@
The usual argument for the existence of WC at a low filling is grounded in approximate energetics calculations and comparisons between WS and FQHL phases using trial wavefunctions, establishing the stability of the WC phase at some low filling factor (typically <1/5<1/5) [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. The serious problems in such numerical calculations are: (1) exclusion of disorder completely; (2) limitations of the trial wavefunction choices for both the WC phase and the FQHL phase. Thus, their reliability in determining $\nu_{c}$ is highly suspect, because the choice of wavefunctions leads to a bias in the energetic comparison, and more importantly, the neglect of disorder makes any application to the emergence of the disordered AS phase in Ref. 1 questionable. In addition, Ref. 1 directly refutes the core argument that the WC phase is stabilized below some critical filling since Ref. 1 clearly observes a WC everywhere except for $\nu<0.093$. Thus, the physics of the crossover to AS is surely not about two competing interacting phases (WC versus FQHL) which applies only for $\nu\approx 1/3$ (but not at $\nu\approx 0.093$ where there is no competing nearby FQHL phase in the STM images), but about the role of disorder in creating the AS phase. The experimentally observed strongly insulating phase at low and nonuniversal filling factors (<1/5<1/5) necessarily requires the existence of disorder, and hence the neglect of disorder in the energetic comparison makes the applicability of the energetic comparison to the experimental transport properties questionable. In addition, Ref. 1 shows that the WC does not become amorphous until the filling factor decreases to a very low value although the usual expectation is that any disorder should pin the WC rendering it insulating at sufficiently weak current and low temperature.
@@
In the current work, we take a radically alternative approach considering only disorder and neglecting all interaction effects to determine where the disorder-induced crossover to the AS phase should occur in the absence of any interaction. This is complementary to the standard approaches to the problem [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 8, 9] where disorder is completely ignored taking into account only interaction with the assumption that any energetic transition to the WC immediately creates a strongly localized insulator. The reality is most likely somewhere in between, but obtaining a complete high-field phase diagram in the presence of both disorder and interaction is one of the most difficult tasks in physics. Our theory below for determining the critical filling factor $\nu_{c}$ is mostly qualitative, at best semi-quantitative, involving drastic approximations and physical arguments since the microscopic problem including both disorder and interaction on an equal footing allowing for the emergence of all three phases (FQHL, WC, AS) observed in Ref. 1 is intractable. Since our focus is the AS phase, we mostly discuss the issue of strong localization in the LLL in the presence of disorder, justifying our findings a posteriori showing that it provides a meaningful explanation for the experimental results on $\nu_{c}$. We emphasize that we use Ref. 1 only as an inspiration and a motivation for our work, and we do not by any means are providing a quantitative theory for the observations of Ref. 1.
@@ II Determination of $\nu_{c}$ and heuristics
As emphasized above, determining the critical filling for the transition to the AS phase from first principles (e.g., QMC based on a microscopic Hamiltonian including both disorder and electron-electron interactions with LL mixing, allowing for all possible phases: FQHL, WC, AS) is essentially impossible since the problem involves strong disorder physics in a non-perturbative interacting problem in the LLL. In fact, just including disorder in the physics of FQHL is an extremely hard problem, which has rarely been attempted (always in very small systems using simplified Hamiltonians) [28, 29, 30]. Using energetics arguments in determining Anderson localization is not a meaningful theoretical approach. Instead, we use several heuristic approaches, which should give a disorder-induced transition at a threshold value of the filling factor, and we discuss their relevance and applicability both to the transition into the AS phase in general and to the specific experimental observations of Ref. 1. Our approach is motivated by our earlier work on the integer quantum Hall effect (IQHE) where we studied in depth the role of disorder, temperature, and LL filling on IQHE in noninteracting 2D systems [31]. We now adapt this approach to the LLL to comment on the localization physics at low filling.
@@
First, we discuss the noninteracting situation in the LLL, which has been studied in great depth recently [31], where all states are localized for $\nu<1/2$ and the Hall conductivity is zero for $\nu<1/2$ with no IQHE. The localization threshold floats up with increasing disorder as shown in the non-interacting phase diagram in Fig. 1 taken from our recent work [31]. Due to the floating-up of extended states, the width of the first $\rho_{xy}=h/e^{2}$ plateau shrinks more rapidly with disorder than that of the second plateau at $\rho_{xy}=2h/e^{2}$, as illustrated in Fig. 1. Consequently, at finite temperature the first plateau can be completely masked by the neighboring broadened extended state, rendering it unobservable in transport measurements. This is totally consistent with experiments—in strongly disordered 2D systems one never observes the basic $\rho_{xy}=h/e^{2}$ quantization because the LLL is completely localized with no extended states. In such a situation, of course, neither the fractional quantum Hall effect (FQHE) nor the WC can manifest itself in the LLL since such interaction-induced phases are much more fragile than the non-interacting IQHE.
@@
Note that this is not an energetic argument, but a more nuanced heuristic involving disorder-induced LL broadening, which, if it becomes comparable to the LL separation, i.e., the cyclotron energy $\hbar\omega_{c}$, there is no IQHE, FQHE, WC—everything in the LLL is then AS with $\nu_{c}>1$. This was in fact the situation for the original discovery of the IQHE by von Klitzing [32], who did not see any LLL IQHE with a quantized Hall resistance of $h/e^{2}$ since the disorder was far too strong in the relevant Si MOS 2D samples he studied. Of course, the original IQHE experiment could not possibly discover any FQHE because of the strong disorder with $\nu_{c}>1$. The FQHE and the associated FQHL evolved later at $\nu=1/3$ with the improvement in samples associated with the switch from dirtier 2D Si MOS samples (with mobility $\mu\sim 10^{3}$ cm2/Vs) to cleaner 2D GaAs samples (with $\mu\sim 10^{5}$ cm2/Vs) [33]. In fact, with this improvement in the quality because of the disorder suppression, $\nu_{c}$ was ostensibly pushed below 1/3 (since 1/3 FQHE was observed), and consequently, the IQHE with $\rho_{xy}=e^{2}/h$ quantization associated with the LLL manifested in this FQHE experiment too [33]. In fact, the manifestation of the $h/e^{2}$ LLL IQHE is a necessary (but not sufficient) condition for the manifestation of any LLL FQHE because $\nu_{c}<1$ is essential for FQHL to form in the LLL. If the whole LLL is a disorder-induced strongly localized AS, there is no QHE (or WC) at all.
@@
We now develop this line of reasoning into a quantitative formula. The basic idea is that the disorder-induced LLL broadening must be smaller than the effective chemical potential (or the Fermi energy) in the LLL for any interaction effects to manifest. If not, the system is an AS. Therefore, by equating the disorder broadening of the LLL to the effective LLL chemical potential, we get the condition defining the critical filling factor for strong localization. This leads to the following equation ($\hbar=1$):
$\Gamma_{\text{LLL}}=\nu_{c}\omega_{c}.$ (1)
@@
Here, $l_{B}=(\hbar c/eB)^{1/2}$ is the magnetic length (or the cyclotron radius in the LLL), and $\tau$ the zero-field disorder scattering time.. $E_{N}=(N+1/2)\hbar\omega_{c}$, $N=0,1,2,3,\dots$, is the energy of the $N$-th LL (with $N=0$ being the LLL). Equation (3) can be intuitively understood as follows. By forming LLs in strong $B$ field, the DOS $D(E)$ of a LL with a width $\Gamma$ increases compared to the parent 2D DOS $D_{0}(E)$ at $B=0$, following $D=D_{0}\hbar\omega_{c}/\Gamma$. Meanwhile, because the DOS increases, the scattering also becomes stronger, and the disorder broadening $\Gamma$ at $B\neq 0$ increases compared to the disorder broadening $\Gamma_{0}=\hbar/\tau$ at $B=0$, following $\Gamma=\Gamma_{0}D/D_{0}$. By solving the two equations $D=D_{0}\hbar\omega_{c}/\Gamma$ and $\Gamma=\Gamma_{0}D/D_{0}$ self-consistently, we obtain $\Gamma^{2}=\hbar\omega_{c}\Gamma_{0}$, which is the same as Eq. (3) up to a numerical factor. In Eqs. (2) and (3), the magnetic field is held fixed, so that $\hbar\omega_{c}$ remains constant, and the filling factor is tuned solely by increasing the carrier density—precisely the same strategy used to control $\nu$ in Ref. 1. The quantity $\tau$ in Eq. (3) is simply the corresponding zero-field disorder scattering time, which, in principle, should be the ‘quantum’ scattering time (also called, the single-particle scattering time), and not the transport relaxation time [35]. The approximations of 1) restricting to the LLL, 2) neglect of LL mixing, and 3) using short-range disorder together considerably simplify the problem to an analytical formula for $\nu_{c}$ which is given by combining Eqs. (1) and (3):
@@
This is again smaller than the experimental $\nu_{c}\sim 0.2$ where the original FQHE sample goes into an insulating phase. Comparing with other high-mobility GaAs samples, we find that the discrepancy between our estimated $\nu_{c}$ and the experimental filling factor, where the system enters into the strong insulator phase in transport, persists in most samples with our estimated $\nu_{c}$ always being typically smaller than the experimental filling factor for the LLL to become insulating. One possible explanation for this discrepancy is that the GaAs samples are modulation-doped, and hence the transport broadening could be orders of magnitude smaller than the quantum broadening since most of the scattering by the remote dopants is forward scattering which does not contribute to transport broadening [35]. It would be quite interesting in this context to measure both the transport and the single particle lifetime in a 2D GaAs sample to see if the discrepancy between theory and experiment is resolved by the enhancement of the quantum broadening over the transport broadening. It is encouraging that the theoretical $\nu_{c}$ is a lower bound since the experimental $\nu_{c}$ always seems larger than the theoretical estimate.
@@
There are alternate heuristic ways to derive analytical expressions for $\nu_{c}$ emphasizing alternate physics in contrast to our emphasis above on the simple equality between the chemical potential and disorder broadening as the criterion for the AS threshold. For example, the above considerations focus solely on the role of disorder by itself in inducing strong localization of the electrons. However, interaction obviously induces a WC phase which is apparent in Fig. 2 of Ref. 1 for $0.2<\nu<0.31$. The question naturally arises what happens to this WC phase in the presence of disorder, which is obviously present in the sample for low enough carrier density, as can be seen for the amorphous structure at $\nu\approx 0.093$. The usual argument is that the WC is pinned by impurities the moment there is any disorder, which may not be showing up at higher values of filling where the carrier density is still high. This may seem problematic because a pinned WC should be an insulator, and transport experiments in high quality GaAs samples do not observe a strongly insulating phase for $\nu>0.2$, which then becomes incompatible with the existence of WC for all $\nu$ except for $\nu\approx 0.334$ in Fig. 2 of Ref. 1. On the other hand, transport experiments on GaAs hole systems reported in Ref. 17 observe insulating behavior extending up to $\nu=1/3$, qualitatively similar to what is seen in BLG. GaAs holes have a larger effective mass and therefore stronger LL mixing than GaAs electrons, while BLG has a much smaller dielectric constant ($\kappa\approx 4$ in hBN, compared to $\kappa\approx 13$ in GaAs), which also enhances LL mixing. Taken together, these observations suggest that stronger LL mixing can shift the critical filling $\nu_{c}$ to higher values. Nevertheless, how $\nu_{c}$ behaves in the presence of both disorder and interaction with LL mixing remains an open question. There is obviously no easy or sharp answer to this question as the disorder-interaction interplay is complex and highly sample dependent. In addition, STM experiments are local whereas transport experiments are global; still the existence of WC everywhere in Ref. 1 poses a challenge to the experimentally observed strongly insulating phase only for very small $\nu$ where Ref. 1 observes an amorphous phase.
@@
We emphasize here that our formula for disorder and field (or density)-dependent $\nu_{c}$ does not incorporate interaction effects, and as such, it is independent of the competing phase being FQHL or WC (or something else). The theory asserts that once $\nu<\nu_{c}$ the system is a randomly spatially localized insulator preempting any other phase existing at that $\nu$ value in the absence of disorder. So, one clear prediction based on our theory is that a gated BLG sample, which is roughly 10 times more disordered, would have a $\nu_{c}$ which is 10 times higher if only the density is being tuned to vary the filling keeping the magnetic field constant. This implies that a sample with 10 times the disorder of Ref. 1 would make a transition into the AS phase at $\nu_{c}\sim 0.93$ if the same magnetic field is being used in the experiment. Therefore, a much dirtier (10 times) sample would manifest no WC or FQHL phase in the LLL, it will only manifest a spatially random quenched electronic AS phase. It is possible that our theory actually becomes more accurate with increasing disorder in which case the transition would happen at 10 times our theoretically estimated $\nu_{c}$ for the current sample of Ref. 1, which would be $\nu_{c}\sim 0.5$. Then, there would be a transition from a WS phase to the AS phase at some rather high filling around $\nu\sim 0.5$, so all FQHL phases will be suppressed by disorder. In either case, this constitutes a striking and readily testable prediction: systematically increasing disorder in the sample should directly reveal whether and how the critical filling shifts.

### v2

@@ I Introduction
First, the system generically displays a hexagonal WC with explicit breaking of the rotational symmetry over a broad range of fillings, and an incompressible FQH liquid appears only in a very narrow window around the odd-denominator fraction such as $\nu=1/3$ Laughlin state (e.g., $\nu\approx 0.334$). Even a slight deviation to $\nu\approx 0.311$ or $0.356$ rapidly restores the WC pattern, consistent with the idea that the $\nu=1/3$ phase is stabilized only within a narrow range close to commensurate odd-denominator fillings. The FQH state slightly lowers its energy compared with the background of a smooth WC phase (existing for all values of $\nu$) through small cusps at these precise odd-denominator $\nu$ (e.g., $1/3$ real space profile as can be seen clearly in Fig. 1 of Ref. 1). Although this is not theoretically unexpected [2, 3, 4], the observation of a generic WC surrounding the $\nu=1/3$ state [1] provides direct verification of this theoretical expectation that FQHL states are stabilized only within a narrow window centered on odd-denominator fillings. This experiment [1] is also the first direct observation of the generic existence of the WC phase in the lowest Landau level of 2D systems.
A new experimental finding in Ref. 1, however, is that, as $\nu$ continues decreasing, eventually at a small $\nu=0.093\approx 1/11$, the electrons cross over to a completely spatially random amorphous solid phase. In the WC, the structure factor $S(q)$ exhibits six well-developed Bragg peaks, reflecting the long-range hexagonal order of the crystal, whereas in the FQHL $S(q)$ is essentially a featureless blob, consistent with the absence of long-range spatial correlations in real space. By contrast, in the amorphous solid phase the real-space images show electrons localized at random, quenched positions, and the corresponding $S(q)$ displays a characteristic hybrid structure: a diffuse central blob accompanied by a surrounding ring. The presence of the ring in $S(q)$ unambiguously identifies the amorphous solid as a solid with a well-defined characteristic length scale, while the central blob signifies liquid-like, short-range spatial correlations. Together, these features establish the amorphous solid as a distinct phase: solid in the sense of electron localization, yet lacking the long-range translational and rotational order characteristic of a crystalline WC. Another impressive recent zero-field density-tuned STM study has reported the observation of a low-density zero-field electronic amorphous phase in highly disordered bilayer MoSe2 samples (no moiré) [5], which has been interpreted as the manifestation of the strongly-localized random disorder-dominated zero-field AS phase in Ref. [6].
@@
The usual argument for the existence of WC at a low filling is grounded in approximate energetics calculations and comparisons between WS and FQHL phases using trial wavefunctions, establishing the stability of the WC phase at some low filling factor (typically <1/5<1/5) [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The serious problems in such numerical calculations are: (1) exclusion of disorder completely; (2) limitations of the trial wavefunction choices for both the WC phase and the FQHL phase. Thus, their reliability in determining $\nu_{c}$ is highly suspect, because the choice of wavefunctions leads to a bias in the energetic comparison, and more importantly, the neglect of disorder makes any application to the emergence of the disordered AS phase in Ref. 1 questionable. In addition, Ref. 1 directly refutes the core argument that the WC phase is stabilized below some critical filling since Ref. 1 clearly observes a WC everywhere except for $\nu<0.093$. Thus, the physics of the crossover to AS is surely not about two competing interacting phases (WC versus FQHL) which applies only for $\nu\approx 1/3$ (but not at $\nu\approx 0.093$ where there is no competing nearby FQHL phase in the STM images), but about the role of disorder in creating the AS phase. The experimentally observed strongly insulating phase at low and nonuniversal filling factors necessarily requires the existence of disorder, and hence the neglect of disorder in the energetic comparison makes the applicability of the energetic comparison to the experimental transport properties questionable. In addition, Ref. 1 shows that the WC does not become amorphous until the filling factor decreases to a very low value. This highlights a crucial distinction: although disorder is expected to pin the WC and cause insulating behavior (at sufficiently weak current and low temperature), a pinned crystal is structurally distinct from a disordered amorphous phase.
@@
In the current work, we take a radically alternative approach considering only disorder and neglecting all interaction effects to determine where the disorder-induced crossover to the AS phase should occur in the absence of any interaction. This is complementary to the standard approaches to the problem [18, 19, 20, 26, 22, 23, 27, 28, 29, 30, 8, 9] where disorder is completely ignored taking into account only interaction with the assumption that any energetic transition to the WC immediately creates a strongly localized insulator. The reality is most likely somewhere in between, but obtaining a complete high-field phase diagram in the presence of both disorder and interaction is one of the most difficult tasks in physics. Our theory below for determining the critical filling factor $\nu_{c}$ is mostly qualitative, at best semi-quantitative, involving drastic approximations and physical arguments since the microscopic problem including both disorder and interaction on an equal footing allowing for the emergence of all three phases (FQHL, WC, AS) observed in Ref. 1 is intractable. Since our focus is the AS phase, we mostly discuss the issue of strong localization in the LLL in the presence of disorder. This approach is justified a posteriori by its ability to provide a meaningful explanation for the experimental results on $\nu_{c}$. We emphasize that we use Ref. 1 only as an inspiration and a motivation for our work, and we do not by any means are providing a quantitative theory for the observations of Ref. 1.
@@ II Determination of $\nu_{c}$ and heuristics
As emphasized above, determining the critical filling for the transition to the AS phase from first principles (e.g., quantum Monte Carlo based on a microscopic Hamiltonian including both disorder and electron-electron interactions with LL mixing, allowing for all possible phases: FQHL, WC, AS) is essentially impossible since the problem involves strong disorder physics in a non-perturbative interacting problem in the LLL. In fact, just including disorder in the physics of FQHL is an extremely hard problem, which has rarely been attempted (always in very small systems using simplified Hamiltonians) [31, 32, 33]. Using energetics arguments in determining Anderson localization is not a meaningful theoretical approach. Instead, we use several heuristic approaches, which should give a disorder-induced transition at a threshold value of the filling factor, and we discuss their relevance and applicability both to the transition into the AS phase in general and to the specific experimental observations of Ref. 1. Our approach is motivated by our earlier work on the integer quantum Hall effect (IQHE) where we studied in depth the role of disorder, temperature, and LL filling on IQHE plateau in noninteracting 2D systems [34]. We now adapt this approach to the LLL to comment on the localization physics at low filling.
@@
First, we discuss the noninteracting situation in the LLL, which has been studied in great depth recently [34], where all states are localized for $\nu<1/2$ and the Hall conductivity is zero for $\nu<1/2$ with no IQHE. The localization threshold floats up with increasing disorder as shown in the non-interacting phase diagram in Fig. 1 taken from our recent work [34]. Due to the floating-up of extended states when disorder is strong $\Gamma>\hbar\omega_{c}$, the width of the first $\rho_{xy}=h/e^{2}$ plateau shrinks more rapidly with disorder than that of the second plateau at $\rho_{xy}=2h/e^{2}$ [34], as illustrated in Fig. 1. Consequently, at finite temperature the first plateau can be completely masked by the neighboring broadened extended state, rendering it unobservable in transport measurements. This is totally consistent with experiments—in strongly disordered 2D systems one never observes the basic $\rho_{xy}=h/e^{2}$ quantization because the LLL is completely localized with no extended states. In such a situation, of course, neither the FQHE nor the WC can manifest itself in the LLL since such interaction-induced phases are much more fragile than the non-interacting IQHE in the presence of strong disorder.
@@
Note that this is not an energetic argument, but a more nuanced heuristic involving disorder-induced LL broadening, which, if it becomes comparable to the LL separation, i.e., the cyclotron energy $\hbar\omega_{c}$, there is no IQHE, FQHE, WC—everything in the LLL is then AS with $\nu_{c}>1$. This was in fact the situation for the original discovery of the IQHE by von Klitzing [35], who did not see any LLL IQHE with a quantized Hall resistance of $h/e^{2}$ since the disorder was far too strong in the relevant Si MOS 2D samples he studied. Of course, the original IQHE experiment could not possibly discover any FQHE because of the strong disorder with $\nu_{c}>1$. The FQHE and the associated FQHL evolved later at $\nu=1/3$ with the improvement in samples associated with the switch from dirtier 2D Si MOS samples (with mobility $\mu\sim 10^{3}$ cm2/Vs) to cleaner 2D GaAs samples (with $\mu\sim 10^{5}$ cm2/Vs) [36]. In fact, with this improvement in the quality because of the disorder suppression, $\nu_{c}$ was pushed below 1/3 (since 1/3 FQHE was observed), and consequently, the IQHE with $\rho_{xy}=e^{2}/h$ quantization associated with the LLL manifested in this FQHE experiment too [36]. The manifestation of the $h/e^{2}$ LLL IQHE is a necessary (but not sufficient) condition for the manifestation of any LLL FQHE because $\nu_{c}<1$ is essential for FQHL to form in the LLL. If the whole LLL is a disorder-induced strongly localized AS, there is no QHE (or WC) at all.
@@
We now develop this line of reasoning into a quantitative formula. The basic idea is that the disorder-induced LLL broadening must be smaller than the effective chemical potential (or the Fermi energy) in the LLL for any interaction effects to manifest. If not, the system is an AS. Therefore, by equating the disorder broadening of the LLL to the effective LLL chemical potential, we get the condition defining the critical filling factor for strong localization. This leads to the following equation:
$\Gamma_{\text{LLL}}=\nu_{c}\hbar\omega_{c}.$ (1)
@@
Here, $l_{B}=(\hbar c/eB)^{1/2}$ is the magnetic length (or the cyclotron radius in the LLL), and $\tau$ the zero-field disorder scattering time.. $E_{N}=(N+1/2)\hbar\omega_{c}$, $N=0,1,2,3,\dots$, is the energy of the $N$-th LL (with $N=0$ being the LLL). Equation (3) can be intuitively understood as follows. By forming LLs in strong $B$ field, the DOS $D(E)$ of a LL with a width $\Gamma$ increases compared to the parent 2D DOS $D_{0}(E)$ at $B=0$, following $D=D_{0}\hbar\omega_{c}/\Gamma$. Meanwhile, because the DOS increases, the scattering also becomes stronger, and the disorder broadening $\Gamma$ at $B\neq 0$ increases compared to the disorder broadening $\Gamma_{0}=\hbar/\tau$ at $B=0$, following $\Gamma=\Gamma_{0}D/D_{0}$. By solving the two equations $D=D_{0}\hbar\omega_{c}/\Gamma$ and $\Gamma=\Gamma_{0}D/D_{0}$ self-consistently, we obtain $\Gamma^{2}=\hbar\omega_{c}\Gamma_{0}$, which is the same as Eq. (3) up to a numerical factor. In Eqs. (2) and (3), the magnetic field is held fixed, so that $\hbar\omega_{c}$ remains constant, and the filling factor is tuned solely by increasing the carrier density—precisely the same strategy used to control $\nu$ in Ref. 1. The quantity $\tau$ in Eq. (3) is simply the corresponding zero-field disorder scattering time, which, in principle, should be the “quantum” scattering time (also called the single-particle scattering time), and not the transport relaxation time [38]. The approximations of 1) restricting to the LLL, 2) neglect of LL mixing, and 3) using short-range disorder together considerably simplify the problem to an analytical formula for $\nu_{c}$ which is given by combining Eqs. (1) and (3):
@@
This is again smaller than the experimental $\nu\sim 0.2$ where the original FQHE GaAs sample goes into an insulating phase. Comparing with other high-mobility GaAs samples, we find that the discrepancy between our estimated $\nu_{c}$ and the experimental filling factor, where the system enters into the strong insulator phase in transport, persists in most samples with our estimated $\nu_{c}$ always being typically smaller than the experimental filling factor for the LLL to become insulating. One possible explanation for this discrepancy is that the experimentally observed insulating phase at $\nu\sim 0.2$ is a pinned WC instead of AS; in this scenario, the predicted AS would emerge at a even lower filling closer to our predicted $\nu_{c}\sim 0.05$ (this scenario is similar to the case in Ref. 1). Alternatively, the discrepancy may arise from the nature of modulation doping in GaAs samples, and hence the transport broadening could be orders of magnitude smaller than the quantum broadening since most of the scattering by the remote dopants is forward scattering which does not contribute to transport broadening [38]. It would be quite interesting in this context to measure both the transport and the single particle lifetime in a 2D GaAs sample to see if the discrepancy between theory and experiment is resolved by the enhancement of the quantum broadening over the transport broadening. It is encouraging that the theoretical $\nu_{c}$ is a lower bound since the experimental $\nu_{c}$ always seems larger than the theoretical estimate.
@@
There are alternative heuristic methods for deriving analytical expressions for $\nu_{c}$ that prioritize different physical mechanisms, contrasting with our focus on the simple equality between the chemical potential and disorder broadening as the criterion for the AS threshold. While our considerations focus primarily on the role of disorder in inducing strong electron localization, electron-electron interactions naturally favor a WC phase, as evidenced in Fig. 2 of Ref. 1 for $0.2<\nu<0.31$. The question naturally arises what happens to this WC phase in the presence of disorder, which is obviously present in the sample for low enough carrier density, as can be seen for the amorphous structure at $\nu\approx 0.093$. The standard expectation is that disorder pins the WC immediately, rendering it insulating. However, transport experiments in high-quality n-GaAs samples do not observe a strongly insulating phase for $\nu>0.2$. This creates an apparent tension with the results in Ref. 1, where a WC is observed across nearly all fillings $0.334>\nu>\nu_{c}\sim 0.093$ except a small range near $\nu\approx 0.334$. In contrast, transport studies on p-GaAs hole systems [17] report insulating behavior extending up to $\nu=1/3$, qualitatively similar to observations in BLG. This discrepancy can be attributed to LL mixing: GaAs holes possess a larger effective mass, and BLG has a significantly lower dielectric constant ($\kappa\approx 4$ in hBN versus $\kappa\approx 13$ in GaAs), both of which enhance LL mixing. Taken together, these observations suggest that enhanced LL mixing shifts the critical filling for the onset of the insulating phase toward higher values. Within this framework, the insulating phase first appears as a pinned WC. As the filling or carrier density decreases, disorder eventually destroys the crystal structure, transitioning the system into an AS at $\nu<\nu_{c}$. This interpretation is consistent with Ref. [1], where the onset of the pinned WC ($\nu\sim 0.4$) occurs at a significantly higher filling than the transition to the AS phase ($\nu\sim\nu_{c}=0.093$). Nevertheless, how $\nu_{c}$ behaves in the presence of both disorder and interaction remains an open question. There is obviously no easy or sharp answer to this question as the disorder-interaction interplay is complex and highly sample dependent.
@@
We emphasize here that our formula for disorder and field (or density)-dependent $\nu_{c}$ does not incorporate interaction effects, and as such, it is independent of the competing phase being FQHL or WC (or something else). The theory asserts that once $\nu<\nu_{c}$ the system is a randomly spatially localized insulator preempting any other phase existing at that $\nu$ value in the absence of disorder. So, one clear prediction based on our theory is that a gated BLG sample, which is roughly 10 times more disordered, would have a $\nu_{c}$ which is 10 times higher if only the density is being tuned to vary the filling keeping the magnetic field constant. This implies that a sample with 10 times the disorder of Ref. 1 would make a transition into the AS phase at $\nu_{c}\sim 0.93$ if the same magnetic field is being used in the experiment. Therefore, a much dirtier (10 times) sample would manifest no WC or FQHL phase in the LLL, it will only manifest a spatially random quenched electronic AS phase. It is possible that our theory actually becomes more accurate with increasing disorder in which case the transition would happen at 10 times our theoretically estimated $\nu_{c}$ for the current sample of Ref. 1, which would be $\nu_{c}\sim 0.5$. Then, there would be a transition from a WC phase to the AS phase at some rather high filling around $\nu\sim 0.5$, so all FQHL phases below $\nu\sim 0.5$ will be suppressed by disorder. This leads to a striking and readily testable prediction: systematically increasing sample disorder should directly reveal this shift in the critical filling.